# PicoPose: Progressive Pixel-to-Pixel Correspondence Learning for Novel Object Pose Estimation

### v1

@@ Abstract
Novel object pose estimation from RGB images presents a significant challenge for zero-shot generalization, as it involves estimating the relative 6D transformation between an RGB observation and a CAD model of an object that was not seen during training. In this paper, we introduce PicoPose, a novel framework designed to tackle this task using a three-stage pixel-to-pixel correspondence learning process. Firstly, PicoPose matches features from the RGB observation with those from rendered object templates, identifying the best-matched template and establishing coarse correspondences. Secondly, PicoPose smooths the correspondences by globally regressing a 2D affine transformation, including in-plane rotation, scale, and 2D translation, from the coarse correspondence map. Thirdly, PicoPose applies the affine transformation to the feature map of the best-matched template and learns correspondence offsets within local regions to achieve fine-grained correspondences. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC. PicoPose achieves state-of-the-art performance on the seven core datasets of the BOP benchmark, demonstrating exceptional generalization to novel objects represented by CAD models or object reference images. Code and models are available at https://github.com/foollh/PicoPose.
Index Terms— Perception for Grasping and Manipulation, Deep Learning for Visual Perception, Machine Learning for Robot Control
I Introduction
Object poses are typically represented by six degrees of freedom (DoFs) parameters, including 3D rotation and translation, to define the transformation from a canonical object space to the camera space. Estimating object poses is highly sought after in real-world applications, such as robotic manipulation [1, 2] and augmented reality [3, 4], and is therefore extensively explored in research.
Early research [5, 6, 7, 8] primarily focused on pose estimation with the same object CAD models for both training and testing phases, but lacked flexibility for the objects unseen during training. Later studies [9, 10, 11, 12, 13] addressed unseen objects within known categories by defining a normalized object coordinate space, but they still struggled with novel categories. With the advancement of foundation models [14, 15, 16], more recent research [17, 18, 19] has increasingly focused on handling entirely new objects to achieve zero-shot 6D object pose estimation, presenting a significant challenge for generalization.
Figure 1: An overview of our PicoPose with a three-stage pixel-to-pixel correspondence learning process for novel object pose estimation from RGB images. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC.
@@ I Introduction
To this end, we introduce a novel framework for progressive pixel-to-pixel correspondence learning, termed as PicoPose, to enable precise pose estimation of novel objects from RGB images. As illustrated in Fig. 1, PicoPose progressively refines the correspondences between RGB observations and templates across three stages, significantly enhancing the accuracy of object poses computed from correspondences.
@@
  • Stage 1: Feature Matching for Coarse Correspondences. In this stage, PicoPose utilizes a visual transformer backbone to capture features for matching the RGB observation and the rendered templates, identifying the best-matched one and obtaining coarse correspondences between the template and the observation.
@@
  • Stage 3: Local Refinement for Fine Correspondences. In this stage, PicoPose applies the affine transformation to the feature map of the best-matched template and employs several offset regression blocks to learn correspondence offsets within local regions, progressively achieving fine-grained correspondences.
@@
To assess the effectiveness of PicoPose, we train the model on the synthetic datasets of ShapeNet-Objects [23] and Google-Scanned-Objects [24] provided by [17], and test it on the seven core datasets of the BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V. The quantitative results on these datasets outperform existing methods, including GigaPose [20] and FoundPose [21], by a significant margin, demonstrating the zero-shot capability of PicoPose. We also successfully apply PicoPose in scenarios where object reference images, rather than object CAD models, are used to represent novel objects. Ablation studies are provided to demonstrate the advantages of individual designs in PicoPose.
@@
In this paper, our key contributions are: (1) the introduction of PicoPose, a novel framework that leverages progressive pixel-to-pixel correspondence learning for pose estimation of novel objects from RGB images; (2) the development of three meticulously designed stages of correspondence learning to improve pose accuracy within PicoPose; and (3) the achievement of state-of-the-art results on the seven core datasets of the BOP benchmark for the RGB-based task.
II Related Work
Figure 2: An illustration of our proposed PicoPose. Given $N$ rendered object templates $\{\mathcal{T}_{i}\}_{i=1}^{N}$ for a target object and a detected RGB crop $\mathcal{I}$ of the observed scene, PicoPose progressively refines the correspondences through three meticulously designed stages to achieve precise pose estimation. In Stage 1, a vision transformer backbone is employed to extract features for template matching and correspondence building, identifying the best-matched template $\mathcal{T}$ and obtaining coarse correspondences between $\mathcal{I}$ and $\mathcal{T}$. In Stage 2, PicoPose encodes the learned correspondences into a correspondence map $\mathcal{A}$, from which it regresses a 2D affine transformation $\mathcal{M}$, including in-plane rotation $\alpha$, scale $s$, and 2D translation $(t_{u},t_{v})$, to smooth the coarse correspondences and filter out outliers. In Stage 3, the transformation $\mathcal{M}$ is applied to the feature map of $\mathcal{T}$, followed by the use of $L$ offset regression blocks to predict coordinate offsets $\Delta\mathcal{P}$ and a certainty map $\mathcal{C}$ within local regions, progressively enhancing the correspondence accuracy. Finally, the object pose is computed using PnP/RANSAC based on these refined correspondences.
II-A Methods Based on Image Matching
@@
To address the generalization challenge, some approaches [26, 27, 17, 20, 19, 28] simplify the task of novel object pose estimation by using an image matching strategy, which involves rendering object templates in various poses and then retrieving the best-matched template to determine the corresponding pose. This strategy is often followed by downstream refinements, as in MegaPose [17] and GenFlow [28]. In contrast, FoundationPose [19] first updates the poses of templates before selecting the best-matched one.
II-B Methods Based on Pixel/Point Matching
@@
This group of methods estimate object poses by establishing correspondences, including 2D-3D correspondences for RGB inputs and 3D-3D correspondences for RGB-D inputs. For instance, OnePose[29] matches the pixel descriptors in object proposals with the point descriptors obtained from Structure from Motion (SfM) to construct the 2D-3D correspondences, and OnePose++[30] further proposes coarse to fine matching to obtain more accurate correspondences. SAM-6D [18] learns 3D-3D correspondences through a two-stage point matching process incorporating background tokens. FoundPose [21] leverages the generalization capabilities of foundation models to extract pixel features and establish 2D-3D correspondences based on feature similarity.
III Method
III-A Overview of PicoPose
The goal of novel object pose estimation from RGB images is to determine the 6D transformation between an RGB observation and a CAD model of an object unseen during training. To address this task, we introduce a novel framework for generalizable pixel-to-pixel correspondence learning, termed as PicoPose, which enables precise estimation of the 6D transformation based on the established correspondences through PnP/RANSAC.
@@ III-A Overview of PicoPose
The architecture of PicoPose is illustrated in Fig. 2. For a given RGB image of a clustered scene and an object CAD model, we begin by rendering object templates from various viewpoints of the CAD model and employing zero-shot segmentation [22] to identify and crop the region containing the target object from the RGB scene. We then resize both the detected crop and the object templates to a fixed size of $H\times W$, representing them as $\mathcal{I}$ and $\{\mathcal{T}_{i}\}_{i=1}^{N}$, respectively, where $N$ is the number of templates. PicoPose utilizes these inputs to search for the best-matched template, denoted as $\mathcal{T}$, and progressively learns the pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$ across three stages, as detailed in Sec. III-B. Each foreground pixel on $\mathcal{T}$ corresponds to a 3D surface point on the CAD model, enabling us to establish pairs of 2D positions on $\mathcal{I}$ and their corresponding 3D points on $\mathcal{T}$, which are utilized for computing the 6D object pose via PnP/RANSAC.
III-B Progressive Pixel-to-Pixel Correspondence Learning of PicoPose
@@
Given the resized detected crop $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and the object templates $\{\mathcal{T}_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$ from various viewpoints of the CAD model, PicoPose employs the ViT-L backbone [31], pretranined by DINOv2 [15], to extract their patch features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{M\times D}$ and $\{\mathcal{F}_{\mathcal{T}_{i}}\in\mathbb{R}^{M\times D}\}_{i=1}^{N}$, respectively, where $D$ is the feature dimension and $M=HW/196$ is the number of patches. PicoPose then identifies the best-matched templates $\mathcal{T}$ and performs three learning stages to build pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$.
@@
Feature similarities between patches of the RGB observation $\mathcal{I}$ and the object templates $\{\mathcal{T}_{i}\}_{i=1}^{N}$, particularly with the best-matched template to $\mathcal{I}$, could establish coarse correspondences.
To enable more effective correspondence establishment, we initially retrieve the best-matched template $\mathcal{T}$ by scoring the degree of similarity between each template $\mathcal{T}_{i}$ and the observation $\mathcal{I}$. For each template $\mathcal{T}_{i}$, we obtain this template matching score $c_{i}$ by averaging the maximum feature cosine similarities of the foreground patches in $\mathcal{I}$ (identified by a prior-step zero-shot segmentation) with the patches in $\mathcal{T}_{i}$, which could be formulated as follows:
$\vspace{-0.1cm}c_{i}=\frac{1}{M^{\prime}}\sum_{j\in\texttt{FG}(\mathcal{F}_{\mathcal{I}})}\max_{k=1,\dots,M}\frac{<\mathbf{f}_{\mathcal{I},j},\mathbf{f}_{\mathcal{T}_{i},k}>}{|\mathbf{f}_{\mathcal{I},j}|\cdot|\mathbf{f}_{\mathcal{T}_{i},k}|},\vspace{-0.1cm}$ (1)
where $\mathbf{f}_{\mathcal{I},j}\in\mathcal{F}_{\mathcal{I}}$ and $\mathbf{f}_{\mathcal{T}_{i},k}\in\mathcal{F}_{\mathcal{T}_{i}}$ are the $j^{th}$ patch features in $\mathcal{I}$ and the $k^{th}$ patch features in $\mathcal{T}_{i}$, respectively, with <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denoting an inner product. $\texttt{FG}(\mathcal{F}_{\mathcal{I}})$ represents the indices of foreground patches in $\mathcal{I}$, and $M^{\prime}$ is the count of foreground patches.
From the set $\{\mathcal{T}_{i}\}_{i=1}^{N}$, the template $\mathcal{T}$ with the highest template matching score is chosen as the best match for establishing pixel-to-pixel correspondences. Furthermore, the feature similarities between $\mathcal{I}$ and $\mathcal{T}$ provide coarse correspondences by determining the most similar patch in $\mathcal{T}$ for each patch in $\mathcal{I}$.
@@
In Stage 1, we exploit feature matching to obtain the coarse and sparse correspondences between $\mathcal{I}$ and $\mathcal{T}$, which, however, often exhibit cluttered distributions with noise and many outliers, as shown in Fig. 1. Therefore, the objective of this stage is to improve the smoothness of these correspondences and filter out the outliers.
@@
where $\alpha$ denotes the in-plane rotation angle, $s$ denotes the relative scale between $\mathcal{I}$ and $\mathcal{T}$, and $(t_{u},t_{v})$ represents the 2D translation of the object centroid in these two images. Applying $\mathcal{M}$ to transform the template $\mathcal{T}$ facilitates pixel alignments with $\mathcal{I}$, thus enabling smooth correspondences. We also note that combining $\mathcal{M}$ with the viewpoint rotation of $\mathcal{T}$ can give the 6D object pose.
As highlighted in Fig. 3, the correspondence map $\mathcal{A}$ between $\mathcal{I}$ and $\mathcal{T}$ can effectively capture the variations in $\alpha$, $s$ and $(t_{u},t_{v})$, thereby encapsulating the essential patterns for learning the affine transformation $\mathcal{M}$ and reducing the associated learning difficulties. Therefore, instead of directly concatenating the patch features $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ for regressing $\mathcal{M}$, we propose a more effective approach by utilizing the coarse correspondences obtained in Stage 1, represented as the correspondence map $\mathcal{A}$ at this stage, to realize the target.
More specifically, we first normalize the feature vectors of $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ to $\overline{\mathcal{F}}_{\mathcal{I}}$ and ¯$\overline{\mathcal{F}}_{\mathcal{T}}$, respectively, and compute the correspondence map $\mathcal{A}$ as $\mathcal{A}=\overline{\mathcal{F}}_{\mathcal{I}}(\overline{\mathcal{F}}_{\mathcal{T}})^{T}\in\mathbb{R}^{M\times M}$, where $M=HW/196$. Subsequently, $\mathcal{A}$ is reshaped to the size of $(H/14)\times(W/14)\times M$ and passed through several stacked convolutions to reduce the spatial dimensions to a global pose vector. Finally, three parallel Multi-layer Perceptrons (MLPs) are applied to the pose vector to learn $(cos(\alpha),sin(\alpha))$, $s$ and $(t_{u},t_{v})$ of $\mathcal{M}$, respectively. Fig. 2 gives the illustration on this stage.
Remark: In GigaPose [20], the affine transformation $\mathcal{M}$ is estimated from each coarse correspondence using a ResNet [32], rather than the visual transformer used in the prior step of template matching, and RANSAC is then applied for selecting the optimal $\mathcal{M}$. In contrast, in our Stage 2, the single estimate of $\mathcal{M}$ can be directly predicted based on the similarities of features extracted by the preceding visual transformer, eliminating the need for RANSAC as a post-processing step and providing auxiliary supervision for training the backbone.
Figure 3: Visualization of correspondence maps between the feature of a point on the RGB observation (marked by a yellow star) and the features of templates with various affine transformations, including in-plane rotation, scaling, and 2D translation.
@@
With the 2D affine transformation $\mathcal{M}$ predicted in Stage 2, we can align the feature map of the best-matched template $\mathcal{T}$ with the RGB observation $\mathcal{I}$, thereby achieving the smooth correspondences. For any position $(u,v)$ on $\mathcal{I}$, the corresponding position $(u^{\prime},v^{\prime})$ on $\mathcal{T}$ could be obtained using $\mathcal{M}$ as follows:
$\begin{bmatrix}u^{\prime}\\ v^{\prime}\end{bmatrix}=\mathcal{M}\begin{bmatrix}u\\ v\\ 1\end{bmatrix}.\vspace{-0.2cm}$ (3)
@@
The deviation between $(u,v)$ and $(u^{\prime},v^{\prime})$ can be interpreted as the commonly known “optical flow” [33]. For all pixels in $\mathcal{I}$, we compute their corresponding positions in $\mathcal{T}$, denoted as $\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$, via Eq. (3) to represent the smooth correspondences. The affine transformation applied to the feature map of $\mathcal{T}$, including rotation, scaling, and translation, is then achieved by using $\mathcal{P}$ as indices to gather features from $\mathcal{T}$, producing a transformed feature map aligned with $\mathcal{I}$. At this stage, we further learn the offsets $\Delta\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$ to update $\mathcal{P}$ to $\mathcal{P}+\Delta\mathcal{P}$, enabling fine-grained correspondence adjustments within local regions.
@@
We realize local refinement of correspondences in a progressively learning manner. Specifically, we first apply Dense Prediction Transformer (DPT) [34] to our backbone, generating $L$ hierarchical feature maps $\{\mathcal{F}_{\mathcal{I}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ and $\{\mathcal{F}_{\mathcal{T}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ of $\mathcal{I}$ and $\mathcal{T}$, respectively, where $H_{l}\times W_{l}$ denotes the spatial size and $D_{l}$ is the number of channels for the $l^{th}$ feature map. We then use $L$ offset regression blocks to iteratively update $\mathcal{P}$.
@@
For the $l^{th}$ offset regression block, the current $\mathcal{P}$ is resized to $H_{l}\times W_{l}\times 2$ and scaled by dividing each 2D position within it by $[H/H_{l}$, $W/W_{l}]$ to ensure spatial consistency. We denote this resized and scaled version as $\mathcal{P}_{l}$, which we use as indices to gather features from $\mathcal{F}_{\mathcal{T}l}$, resulting in the transformed feature map $\mathcal{F}_{\mathcal{T}l}^{\prime}$ to align with $\mathcal{F}_{\mathcal{I}l}$. Additionally, we introduce a third correlation feature map $\mathcal{F}_{\mathcal{C}l}$ via a Correlation Lookup module, introduced in RAFT [35], to explicitly provide correlation degrees and facilitate easier learning of offsets; more details on this module can be found in the RAFT paper [35]. We then concatenate $\mathcal{F}_{\mathcal{I}l}$, $\mathcal{F}_{\mathcal{T}l}^{\prime}$, and $\mathcal{F}_{\mathcal{C}l}$ to form the input for two sequences of stacked convolutions, which are used to regress offsets $\Delta\mathcal{P}_{l}\in\mathbb{R}^{H_{l}\times W_{l}\times 2}$ and certainty map $\mathcal{S}_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$. $\Delta\mathcal{P}_{l}$ is then interpolated to the size of $H\times W\times 2$, scaled by multiplying the 2D coordinates within it by $[H/H_{l},W/W_{l}]$, and added to $\mathcal{P}$ for updating. The certainty map $\Delta\mathcal{S}_{l}$ represents the confidence of the regressed offsets and is upsampled to produce $\mathcal{C}_{l}^{\prime}\in\mathbb{R}^{H\times W}$.

### v2

@@ Abstract
RGB-based novel object pose estimation is critical for rapid deployment in robotic applications, yet zero-shot generalization remains a key challenge. In this paper, we introduce PicoPose, a novel framework designed to tackle this task using a three-stage pixel-to-pixel correspondence learning process. Firstly, PicoPose matches features from the RGB observation with those from rendered object templates, identifying the best-matched template and establishing coarse correspondences. Secondly, PicoPose smooths the correspondences by globally regressing a 2D affine transformation, including in-plane rotation, scale, and 2D translation, from the coarse correspondence map. Thirdly, PicoPose applies the affine transformation to the feature map of the best-matched template and learns correspondence offsets within local regions to achieve fine-grained correspondences. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC. PicoPose achieves state-of-the-art performance on the seven core datasets of the BOP benchmark, demonstrating exceptional generalization to novel objects. Code and trained models are available at https://github.com/foollh/PicoPose.
Keywords: Novel Object Pose Estimation, Robotic Manipulation
1 Introduction
Object poses are typically represented by six degrees of freedom (DoFs) parameters, including 3D rotation and translation, to define the transformation from a canonical object space to the camera space. Estimating object poses is highly sought after in real-world applications, particularly in the context of robotics, where it enables precise manipulation, grasping, and interaction with various objects [1, 2, 3, 4, 5] (see Fig. 1), and is therefore extensively explored in research.
Early research [6, 7, 8, 9] primarily focused on pose estimation with the same object CAD models for both training and testing phases, but lacked flexibility for the objects unseen during training. Later studies [10, 11, 12, 13, 14] addressed unseen objects within known categories by defining a normalized object coordinate space, but struggled with novel categories. With the advancement of foundation models [15, 16, 17], recent research [18, 19, 20] has increasingly focused on handling new objects to achieve zero-shot pose estimation. While this capability enables rapid deployment of robotic systems, it presents significant generalization challenges that remain to be addressed.
Figure 1: An overview of our proposed PicoPose with a three-stage pixel-to-pixel correspondence learning process for novel object pose estimation from RGB images. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC. With zero-shot capability, PicoPose enables rapid deployment across various robotic manipulation systems for unseen objects.
@@ 1 Introduction
To this end, we introduce a novel framework for progressive pixel-to-pixel correspondence learning, termed as PicoPose, to enable precise pose estimation of novel objects from RGB images. As illustrated in Fig. 1, PicoPose progressively refines the correspondences between RGB observations and templates across three stages, significantly enhancing the accuracy of object poses. With zero-shot capability, PicoPose enables rapid deployment across various robotic manipulation systems.
@@
  • Stage 1: Feature Matching for Coarse Correspondences. In this stage, PicoPose utilizes visual transformers to capture features for matching the RGB observation and the rendered templates, identifying the best-matched one and obtaining coarse correspondences.
@@
  • Stage 3: Local Refinement for Fine Correspondences. In this stage, PicoPose applies the affine transformation to the feature map of the best-matched template and employs several offset regression blocks to learn correspondence offsets within local regions.
@@
We train PicoPose on the synthetic datasets of ShapeNet-Objects [24] and Google-Scanned-Objects [25], and test it on seven BOP datasets [26]. The quantitative results on these datasets outperform other existing methods, demonstrating the zero-shot capability of PicoPose. We also apply PicoPose in scenarios where object reference images are used to represent novel objects.
@@
In this paper, our key contributions are: (1) the introduction of PicoPose, a novel framework that leverages progressive pixel-to-pixel correspondence learning for pose estimation of novel objects from RGB images; (2) the development of three designed stages of correspondence learning to improve pose accuracy within PicoPose; and (3) the achievement of state-of-the-art results on the seven core datasets of the BOP benchmark for the RGB-based task, especially without refinement.
2 Related Work
@@
Methods Based on Image Matching. To address the challenge of generalization, some approaches [27, 28, 18, 21, 20, 29] simplify the task of novel object pose estimation by using an image matching strategy, which involves rendering object templates in various poses and then retrieving the best-matched template to determine the corresponding pose. This strategy is often followed by downstream refinements, as in MegaPose [18] and GenFlow [29]. In contrast, FoundationPose [20] first updates the poses of templates before selecting the best-matched one.
@@
Methods Based on Pixel/Point Matching. This group of methods estimates object poses by establishing correspondences, including 2D-3D correspondences for RGB inputs and 3D-3D correspondences for RGB-D inputs. For instance, OnePose [30] matches the pixel descriptors in object proposals with the point descriptors obtained from Structure from Motion (SfM) to construct the 2D-3D correspondences, and OnePose++ [31] further proposes coarse to fine matching to obtain more accurate correspondences. SAM-6D [19] learns 3D-3D correspondences through a two-stage point matching process incorporating background tokens. FoundPose [22] leverages the generalization capabilities of foundation models to extract pixel features and establish 2D-3D correspondences.
3 Method
3.1 Overview of PicoPose
@@
The goal of novel object pose estimation from RGB images is to determine the 6D transformation between an RGB observation and a CAD model of an object unseen during training. To address this task, we introduce a novel framework for generalizable pixel-to-pixel correspondence learning, termed as PicoPose. The architecture of PicoPose is illustrated in Fig. 2. For a given RGB image of a clustered scene and an object CAD model, we begin by rendering object templates from various viewpoints of the CAD model and employing zero-shot segmentation [23] to identify and crop the region containing the target object from the RGB scene. We then resize both the detected crop and the object templates to a fixed size of $H\times W$, representing them as $\mathcal{I}$ and $\{\mathcal{T}_{i}\}_{i=1}^{K}$, respectively, where $K$ is the number of templates. PicoPose utilizes these inputs to search for the best-matched template, denoted as $\mathcal{T}$, and progressively learns the pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$ across three stages, as detailed in Sec. 3.2. Each foreground pixel on $\mathcal{T}$ corresponds to a 3D surface point on the CAD model, enabling us to establish pairs of 2D positions on $\mathcal{I}$ and their corresponding 3D points on $\mathcal{T}$, which are utilized for computing the 6D object pose via PnP/RANSAC.
Figure 2: An illustration of our proposed PicoPose.
3.2 Progressive Pixel-to-Pixel Correspondence Learning of PicoPose
@@
Given the resized detected crop $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and the object templates $\{\mathcal{T}_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{K}$ , PicoPose employs the ViT-L backbone [32], pretranined by DINOv2 [16], to extract their patch features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{N\times D}$ and $\{\mathcal{F}_{\mathcal{T}_{i}}\in\mathbb{R}^{N\times D}\}_{i=1}^{K}$, respectively, where $D$ is the feature dimension and $N$ is the number of patches. PicoPose then identifies the best-matched templates $\mathcal{T}$ and performs three learning stages to build pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$.
@@
Feature similarities between patches of the RGB observation $\mathcal{I}$ and the object templates $\{\mathcal{T}_{i}\}_{i=1}^{K}$, particularly with the best-matched template to $\mathcal{I}$, could establish coarse correspondences. To make it more effective, we initially retrieve the best-matched template $\mathcal{T}$ by scoring the degree of similarity between each template $\mathcal{T}_{i}$ and the observation $\mathcal{I}$. For each template $\mathcal{T}_{i}$, we obtain this template matching score $c_{i}$ by averaging the maximum feature cosine similarities of the foreground patches in $\mathcal{I}$ (identified by a prior-step zero-shot segmentation) with the patches in $\mathcal{T}_{i}$ as follows:
$c_{i}=\frac{1}{N^{\prime}}\sum_{j\in\texttt{FG}(\mathcal{F}_{\mathcal{I}})}\max_{k=1,\dots,N}\frac{<\mathbf{f}_{\mathcal{I},j},\mathbf{f}_{\mathcal{T}_{i},k}>}{|\mathbf{f}_{\mathcal{I},j}|\cdot|\mathbf{f}_{\mathcal{T}_{i},k}|},$ (1)
where $\mathbf{f}_{\mathcal{I},j}\in\mathcal{F}_{\mathcal{I}}$ and $\mathbf{f}_{\mathcal{T}_{i},k}\in\mathcal{F}_{\mathcal{T}_{i}}$ are the $j^{th}$ patch features in $\mathcal{I}$ and the $k^{th}$ patch features in $\mathcal{T}_{i}$, respectively, with $\cdot$$\cdot$> denoting an inner product. $\texttt{FG}(\mathcal{F}_{\mathcal{I}})$ represents the indices of foreground patches in $\mathcal{I}$, and $N^{\prime}$ is the count of foreground patches. From $\{\mathcal{T}_{i}\}_{i=1}^{K}$, the template $\mathcal{T}$ with the highest score is chosen as the best match. Furthermore, the feature similarities between $\mathcal{I}$ and $\mathcal{T}$ give coarse correspondences by determining the most similar patch in $\mathcal{T}$ for each patch in $\mathcal{I}$.
@@
In Stage 1, we exploit feature matching to obtain the coarse and sparse correspondences, which, however, often exhibit cluttered distributions with noise and many outliers. Therefore, the objective of this stage is to improve the smoothness of these correspondences and filter out the outliers.
@@
where $\alpha$ denotes the in-plane rotation angle, $s$ denotes the relative scale between $\mathcal{I}$ and $\mathcal{T}$, and $(t_{u},t_{v})$ represents the 2D translation of the object centroid in these two images. Applying $\mathcal{M}$ to transform the template $\mathcal{T}$ facilitates pixel alignments with $\mathcal{I}$, thus enabling smooth correspondences.
As highlighted in Fig. 3, the correspondence map $\mathcal{A}$ between $\mathcal{I}$ and $\mathcal{T}$ can effectively capture the variations in $\alpha$, $s$, and $(t_{u},t_{v})$, thereby encapsulating the essential patterns for learning $\mathcal{M}$. Therefore, instead of directly concatenating the patch features $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ for regressing $\mathcal{M}$, we propose a more effective approach by utilizing the coarse correspondences obtained in Stage 1, represented as the correspondence map $\mathcal{A}$ at this stage, to realize the target. More specifically, we first normalize the feature vectors of $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ to $\overline{\mathcal{F}}_{\mathcal{I}}$ and $\overline{\mathcal{F}}_{\mathcal{T}}$, respectively, and compute the correspondence map $\mathcal{A}$ as $\mathcal{A}=\overline{\mathcal{F}}_{\mathcal{I}}(\overline{\mathcal{F}}_{\mathcal{T}})^{T}\in\mathbb{R}^{N\times N}$, where $N=HW/196$. Subsequently, $\mathcal{A}$ is reshaped to the size of $(H/14)\times(W/14)\times N$ and passed through several stacked convolutions to reduce the spatial dimensions to a global pose vector. Finally, three parallel Multi-layer Perceptrons (MLPs) are applied to the pose vector to learn $(cos(\alpha),sin(\alpha))$, $s$ and $(t_{u},t_{v})$ of $\mathcal{M}$, respectively.
Figure 3: Visualization of correspondence maps between the feature of a point on the RGB observation (marked by a yellow star) and the features of templates with various affine transformations.
@@
With the transformation $\mathcal{M}$ predicted in Stage 2, we can align the feature map of the best-matched template $\mathcal{T}$ with the RGB observation $\mathcal{I}$, thereby achieving the smooth correspondences. For any position $(u,v)$ on $\mathcal{I}$, the corresponding position $(u^{\prime},v^{\prime})$ on $\mathcal{T}$ could be obtained using $\mathcal{M}$ as follows:
$\begin{bmatrix}u^{\prime}\\ v^{\prime}\end{bmatrix}=\mathcal{M}\begin{bmatrix}u\\ v\\ 1\end{bmatrix}$. (3)
@@
The deviation between $(u,v)$ and $(u^{\prime},v^{\prime})$ can be interpreted as the commonly known “optical flow” [33]. For all pixels in $\mathcal{I}$, we compute their corresponding positions in $\mathcal{T}$, denoted as $\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$, via Eq. (3) to represent the smooth correspondences. The affine transformation applied to the feature map of $\mathcal{T}$ is then achieved by using $\mathcal{P}$ as indices to gather features from $\mathcal{T}$, producing a transformed feature map aligned with $\mathcal{I}$. At this stage, we further learn the offsets $\Delta\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$ to update $\mathcal{P}$ to $\mathcal{P}+\Delta\mathcal{P}$, enabling fine-grained correspondence adjustments within local regions.
@@
We realize local refinement of correspondences in a progressive learning manner. Specifically, we first apply Dense Prediction Transformer (DPT) [34] to our backbone, generating $L$ hierarchical feature maps $\{\mathcal{F}_{\mathcal{I}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ and $\{\mathcal{F}_{\mathcal{T}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ of $\mathcal{I}$ and $\mathcal{T}$, respectively, where $H_{l}\times W_{l}$ denotes the spatial size and $D_{l}$ is the number of channels for the $l^{th}$ feature map. We then use $L$ offset regression blocks to iteratively update $\mathcal{P}$.
@@
For the $l^{th}$ offset regression block, the current $\mathcal{P}$ is resized to $H_{l}\times W_{l}\times 2$ and scaled by dividing each 2D position within it by $H/H_{l}$, $W/W_{l}]$ to ensure spatial consistency. We denote this resized and scaled version as $\mathcal{P}_{l}$, which we use as indices to gather features from $\mathcal{F}_{\mathcal{T}l}$, resulting in the transformed feature map $\mathcal{F}_{\mathcal{T}l}^{\prime}$ to align with $\mathcal{F}_{\mathcal{I}l}$. Additionally, we introduce a third correlation feature map $\mathcal{F}_{\mathcal{C}l}$ via a Correlation Lookup module, introduced in RAFT [35], to explicitly provide correlation degrees and facilitate easier learning of offsets; more details on this module can be found in the RAFT paper [35]. We then concatenate $\mathcal{F}_{\mathcal{I}l}$, $\mathcal{F}_{\mathcal{T}l}^{\prime}$, and $\mathcal{F}_{\mathcal{C}l}$ to form the input for two sequences of stacked convolutions, which are used to regress offsets $\Delta\mathcal{P}_{l}\in\mathbb{R}^{H_{l}\times W_{l}\times 2}$ and certainty map $\mathcal{S}_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$. $\Delta\mathcal{P}_{l}$ is then interpolated to the size of $H\times W\times 2$, scaled by multiplying the 2D coordinates within it by $H/H_{l},W/W_{l}]$, and added to $\mathcal{P}$ for updating. The certainty map $\Delta\mathcal{S}_{l}$ represents the confidence of the regressed offsets and is upsampled to generate $\mathcal{C}_{l}^{\prime}\in\mathbb{R}^{H\times W}$.